# Attractor Stability of Boolean networks under noise

### v1

@@ IV.2 Dominance and stationary distribution of attractors
Figure 4: Comparison between the empirically measured time fractions spent in attractors, $t_{\alpha}/t_{\text{tot}}$, and the principal eigenvector components $v_{\alpha}$ of the transition matrix is shown for random regular networks with $N=20$ and $k=2,5$. Panels (a,c) show comparisons with eigenvector predictions, while panels (b,d) show comparisons with static basin sizes.
@@
The transition matrix provides a framework for identifying the dynamical importance of attractors under noise. As discussed in Sec. IIB, its principal eigenvector $\vec{v}$ offers a theoretical estimate of the fraction of time spent in each attractor. To validate this prediction, we measure the average duration that the system spends in different attractors of Boolean networks under noise. For each simulation, we track the transitions between attractors and record the time $t_{\alpha}$ spent in each attractor $\alpha$. Specifically, we measure the time fractions $t_{\alpha}/t_{\text{tot}}$, where $t_{\text{tot}}$ is the total simulation time.
@@
In Fig. 4, we compare the components $v_{\alpha}$ of principal eigenvector with the numerically measured time fractions. Figures 4(a,c) show that the eigenvector accurately predicts the stationary distribution, while Figs. 4(b,d) demonstrate that basin size provides a less accurate estimate of the empirical distribution. The strong agreement between $v_{\alpha}$ and the measured values $t_{\alpha}/t_{\text{tot}}$ confirms that the system’s stationary distribution is well predicted by the leading eigenvector of the transition matrix. These results suggest that the steady state of Boolean networks are governed by noise-induced transition probabilities rather than static basin structures.
Additionally, this correspondence implies that attractor dynamics can be interpreted as a random walk on the attractor transition network, with transitions governed by the probabilities $m_{\alpha\beta}$. This is conceptually related to the PageRank algorithm [37] or eigenvector centrality [38, 39, 40, 41], in which the stationary distribution of a Markov chain determines node importance. In our framework, attractors play the role of nodes, and their importance is reflected in the time that the system spends in each under noise predicted theoretically by the leading eigenvector. Thus, this approach provides an efficient way to identify dominant attractors and to predict the asymptotic behavior of Boolean networks under stochastic perturbations.
@@ IV.3 Entropy of attractor distributions
We further study how broadly the system explores the attractor space under noise. To this end, we compute the normalized Shannon entropy $H$ of the principal eigenvector $\vec{v}$ of the transition matrix, defined as
@@
where $v_{\alpha}$ represents the steady-state probability that the system occupies attractor $\alpha$ under noise, and $H_{\max}=\log_{2}n_{\text{att}}$ is the maximum entropy possible for a given number of attractors. The normalization ensures that $H\in[0,1]$. High entropy indicates that the system frequently visits many attractors with similar probabilities, reflecting broad exploration and dynamical diversity. In contrast, low entropy implies strong localization around a few dominant attractors. Figure 5(a) shows that the Shannon entropy $H$ as a function of $s$ for various $k$. This result shows a shift from localized to distributed dynamics as $s$ varies.
To assess how this behavior compares with the static structure of the state space, we compute the entropy difference $\Delta H$ between the stationary distribution and the basin size distribution:
@@
where $b_{\alpha}$ is the normalized basin size of attractor $\alpha$. Positive values of $\Delta H$ indicate that the system explores a narrower subset of attractors under local stochastic dynamics than would be expected based on basin sizes alone. Figure 5(b) shows that $\Delta H$ is generally positive, particularly at low sensitivity, showing that noise-driven dynamics remain localized despite the existence of other attractors. As $s$ increases, $\Delta H$ gradually decreases and hovers near zero, suggesting convergence toward the global randomization predicted by basin sizes.
To better understand this localization, we examine transition asymmetries between attractors. Figure 5(c) shows the ratio of transition probabilities between attractors of different basin sizes. We compute the ratio $\frac{m_{\alpha\beta}}{b_{\alpha}}$, where $m_{\alpha\beta}$ is the probability of transitioning from attractor $\alpha$ to $\beta$, and $b_{\alpha}$ is the basin size of the source attractor. We then compare this value for transitions into larger basins ($R_{>}$) versus smaller ones ($R_{<}$), using the ratio $R_{>}/R_{<}$ to quantify the directionality of transitions. Our results indicate that for $s<1.5$, transitions tend to favor smaller basins, i.e., $R_{>}/R_{<}<1$, suggesting a bias in the system’s exploration. That is, attractors with larger basin sizes become more dominant in this regime, leading to localization around a limited set of attractors. At higher sensitivities, this bias diminishes and the ratio approaches $1$, implying that the transition dynamics become more random and basin-size neutral, consistent with global randomization.
@@ IV.4 Application to real-world Boolean networks
Finally, we apply our framework to real-world Boolean networks obtained from the Cell Collective database [42, 43]. These datasets include biological systems such as gene regulatory networks, signaling pathways, and developmental processes. We found that the trace of the transition matrix, $\mathrm{Tr}(M)$, exceeded unity consistently across all tested networks as shown in Fig. 6(a). This indicates that, similar to random Boolean networks, real-world Boolean networks exhibit enhanced attractor stability under local stochastic perturbations. This result suggests that high stability to noise is not an artifact of synthetic models but a robust feature observed in biological Boolean systems.
We further compared the stationary distributions predicted by the principal eigenvectors with the basin size distributions. In Fig. 6(b), each data point corresponds to an individual attractor from 17 different real-world Boolean networks, where the horizontal axis represents the normalized basin size and the vertical axis indicates the corresponding component of the principal eigenvector. We found that basin sizes deviate from eigenvector components, indicating that static basin structures are insufficient to predict the long-term dynamics under noise. Since noise is inherent in real-world systems, our framework based on transition dynamics provides a reliable prediction on the behavior of real-world Boolean networks.
@@ V Discussion
In this study, we developed a framework to quantify the stability and transition dynamics of attractors in Boolean networks under noise. By constructing and analyzing transition matrices, we showed that attractors are more stable against local noise than predicted by static basin sizes. We also found that basin sizes alone are insufficient to predict long-term dynamics under noise, and that transition dynamics allows us a more reliable predictions. We further demonstrated that real-world biological networks exhibit similar resilience to noise, suggesting that this can be a general feature of biological systems.
Our study provides an alternative perspective on attractors of Boolean networks, showing the importance of dynamical structure and stochastic effects in noisy environments. These findings suggest that noise can induce structured yet nontrivial transitions between attractors, and that Boolean networks exhibit an intrinsic resilience. It remains important to assess the generality of these findings, especially in large-scale networks, as a key direction for future work. In addition, future research could explore how specific topological features, such as modularity [18, 44], redundancy [17, 45], and/or the presence of canalizing functions [46, 14, 15], contribute to this stability. Extending the framework to include asynchronous updates [13], multi-node perturbations [35], or correlated noise [47] may provide further insight into biological function.
@@ Acknowledgements.
This research was supported in part by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. 2020R1I1A3068803) and by Global - Learning & Academic research institution for Master’s $\cdot$ PhD students, and Postdocs (LAMP) Program of the National Research Foundation of Korea (NRF) grant funded by the Ministry of Education (No. RS-2024-00445180). This work was also supported by a funding for the academic research program of Chungbuk National University in 2025. The research of R.L. was partially supported by the grants NSF DMS-2424635, NIH R01 AI135128, and NIH R01 HL169974-01.

### v2

@@ IV.1 Dominance and stationary distribution of attractors
Figure 2: Comparison between the empirically measured time fractions spent in attractors, $t_{\alpha}/t_{\text{tot}}$, and the principal eigenvector components $v_{\alpha}$ of the transition matrix is shown for random regular networks with $N=20$ and $k=2,5$. Panels (a,c) show comparisons with eigenvector predictions, while panels (b,d) show comparisons with basin sizes.
@@
The transition matrix provides an efficient way for identifying the dynamical importance of attractors under noise. As discussed in Sec. IIIB, its principal eigenvector $\vec{v}$ offers a theoretical estimate of the fraction of time spent in each attractor. To validate this prediction, we measure the average duration that the system spends in different attractors of Boolean networks under noise. For each simulation, we track the transitions between attractors and record the time $t_{\alpha}$ spent in each attractor $\alpha$. Specifically, we measure the time fractions $t_{\alpha}/t_{\text{tot}}$, where $t_{\text{tot}}$ is the total simulation time. In these simulations, the noise rate is set to $0.001$, and one time step corresponds to an update of all nodes.
In Fig. 2, we compare the components $v_{\alpha}$ of principal eigenvector with the numerically measured time fractions. Figures 2(a,c) show that the eigenvector accurately predicts the stationary distribution, while Figs. 2(b,d) show the relationship between basin sizes and the empirical time fractions. The Pearson correlation between $v_{\alpha}$ and the measured values $t_{\alpha}/t_{\text{tot}}$ exceeds 0.99, indicating the system’s stationary distribution is well predicted by the leading eigenvector of the transition matrix. The small discrepancies arise when noise occurs during transient periods before the system fully relaxes back to an attractor.
Additionally, this correspondence implies that attractor dynamics can be interpreted as a random walk on the attractor transition network, with transitions governed by the probabilities $m_{\alpha\beta}$. This is conceptually related to the PageRank algorithm [41] or eigenvector centrality [42, 43, 44], in which the stationary distribution of a Markov chain determines node importance. In our framework, attractors play the role of nodes, and their importance is reflected in the time that the system spends in each under noise predicted theoretically by the leading eigenvector. Thus, this approach provides an efficient way to identify dominant attractors and to predict the asymptotic behavior of Boolean networks under stochastic perturbations.
IV.2 Stability of attractors under noise
Figure 3: (a) The trace $\mathrm{Tr}(M)$ of transition matrix for random Boolean networks on random regular networks with size $N=20$ and various degrees $k$ with respect to the sensitivity $s$ is shown. (b) The trace normalized by the number of attractors, $\mathrm{Tr}(M)/n_{\mathrm{att}}$ is shown. Inset shows $\mathrm{Tr}(M)/n_{\mathrm{att}}$ as a function of size NN for $(k,s)=(2,0.75)$ for squares and $(k,s)=(4,1.5)$ for triangles.
An important concept in Boolean networks is the stability of attractors, which refers to the system’s ability to return to the same attractor after a perturbation. The stability of an attractor $\alpha$ can be directly quantified by the diagonal element $m_{\alpha\alpha}$ of the transition matrix. To evaluate stability at the network level, we compute the trace of the transition matrix, $\mathrm{Tr}(M)$. A larger trace indicates that attractors in Boolean networks are more stable to noise. For local noise, the trace can be computed from the transition matrix constructed by explicitly evaluating the effects of all local perturbations. For the global randomization in which the system’s state is reset entirely at random, each term $m_{\alpha\alpha}$ becomes $b_{\alpha}/N$ where $b_{\alpha}$ is the basin size. Thus, the trace of the transition matrix equals one by definition, as it results from the normalization of basin sizes. Thus, we obtain $\mathrm{Tr}(M)=1$ for the global noise.
We compare the stability of attractors for local and global noise in Fig. 3(a). The trace $\mathrm{Tr}(M)$ for local noise is consistently greater than unity across various values of the connectivity $k$ and sensitivity $s$, on random regular networks with $N=20$. These observations suggest that Boolean networks have an intrinsic resilience to local perturbations, a feature that has been recognized since the pioneering work in Boolean networks [1]. Our findings also suggest the importance of assessing stability through $\mathrm{Tr}(M)$, which directly reflect the system’s resilience to noise.
In Fig. 3(b), we show the average attractor stability under local noise, defined as $\mathrm{Tr}(M)/n_{\mathrm{att}}$, where $n_{\mathrm{att}}$ is the number of attractors in a given network. This normalization is introduced because the number of attractors varies across different network instances and parameter settings. The results show that average attractor stability tends to decrease as the average sensitivity $s$ increases. The inset of Fig. 3(b) shows the relation between $\mathrm{Tr}(M)/n_{\mathrm{att}}$ and the network size $N$, which appears nearly constant within the range studied. However, because the tested sizes are small, the nature of this size dependence requires further investigation. This behavior is also different from the global-noise case, where the attractor stability becomes $1/n_{\mathrm{att}}$ and is independent of $s$ and $k$.
In the case of local noise, we can also estimate a lower bound for attractor stability, which depends on the network parameters $s$ and $k$. Within the annealed approximation [21], we can estimate the one-step probability $P_{\mathrm{ret}}^{(1)}$ that a single-node perturbation disappears after one update step, i.e., that the system returns exactly to the pre-perturbation state, as
$P_{\mathrm{ret}}^{(1)}\approx\left(1-\frac{s}{k}\right)^{k}.$ (5)
This quantity becomes a lower bound on the attractor stability, because perturbed states may still belong to the same attractor basin even when they do not immediately return to the original state. This expression captures the dependence of the stability on the parameters $s$ and $k$, and it also shows that the bound decreases monotonically as the sensitivity $s$ increases.
@@ IV.3 Entropy of attractor distributions
We further study how broadly the system explores the attractor space. To this end, we compute the normalized Shannon entropy $H$ of the principal eigenvector $\vec{v}$ of the transition matrix, defined as
@@
where $v_{\alpha}$ represents the steady-state probability that the system occupies attractor $\alpha$ under noise, and $H_{\max}=\log_{2}n_{\text{att}}$ is the maximum entropy possible for a given number of attractors. The normalization ensures that $H\in[0,1]$. High entropy indicates that the system frequently visits many attractors with similar probabilities, reflecting broad exploration and dynamical diversity. In contrast, low entropy implies strong localization around a few dominant attractors. Figure 4(a) shows that the Shannon entropy $H$ as a function of $s$ for various $k$.
To assess how this behavior compares with the global randomization, we compute the entropy difference $\Delta H$ between the stationary distribution and the basin size distribution:
@@
where $b_{\alpha}$ is the normalized basin size of attractor $\alpha$. Positive values of $\Delta H$ indicate that the system explores a broader subset of attractors under local stochastic dynamics than would be expected based on the global randomization. Figure 4(b) shows that $\Delta H$ is positive at low sensitivity, showing that attractor dynamics become more spread out than what the global randomization would predict. As $s$ increases, $\Delta H$ gradually decreases and hovers near zero above $s>1.5$, suggesting convergence toward the global randomization.
We examine transition asymmetries between attractors. Figure 4(c) shows the ratio of transition probabilities between attractors of different basin sizes. We compute the ratio $\frac{m_{\alpha\beta}}{b_{\alpha}}$, where $m_{\alpha\beta}$ is the probability of transitioning from attractor $\alpha$ to $\beta$, and $b_{\alpha}$ is the basin size of the source attractor. We then compare this value for transitions into larger basins ($R_{>}$) versus smaller ones ($R_{<}$), using the ratio $R_{>}/R_{<}$ to quantify the directionality of transitions. Our results indicate that for $s<1.5$, transitions tend to favor smaller basins, i.e., $R_{>}/R_{<}<1$. This pattern of transitions enhances the occupation of small-basin attractors relative to their basin sizes. At higher sensitivities, this tendency diminishes and the ratio approaches $1$, implying that the transition dynamics become more random, consistent with global randomization.
@@ IV.4 Application to real-world Boolean networks
Finally, we apply our framework to real-world Boolean networks obtained from the Cell Collective database [45, 46]. These datasets include biological systems such as gene regulatory networks, signaling pathways, and developmental processes. Basic properties of the real-world Boolean networks used in this section are presented in the Table 1. in the Appendix. We found that the trace of the transition matrix, $\mathrm{Tr}(M)$, exceeded unity consistently across all tested networks as shown in Fig. 5(a). This indicates that, similar to random Boolean networks, real-world Boolean networks exhibit enhanced attractor stability under local noise.
We further compared the stationary distributions predicted by the principal eigenvectors with the basin size distributions. In Fig. 5(b), each data point corresponds to an individual attractor from 17 different real-world Boolean networks, where the horizontal axis represents the normalized basin size and the vertical axis indicates the corresponding component of the principal eigenvector. We found that basin sizes deviate from eigenvector, showing that local and global noise produce distinct patterns of attractor dynamics. Some extreme data points with larger basin size and $v=0$ correspond to unused attractors in noisy environments that still possess relatively large basin sizes. These unusual attractors show interesting anomalies and become a potentially direction for future research.
@@ V Discussion
In this study, we proposed a framework for analyzing attractor dynamics in Boolean networks under noise, based on a representation of transitions between attractors. This approach offers insight into key properties of attractor behavior such as long-term dominance and stability. We also compare the local and global noise, showing that they lead to different behaviors in attractor dynamics. To be specific, local noise shows structured and non-trivial transition patterns, whereas global noise yields attractor dynamics largely determined by basin sizes.
Further applications of our method to real-world Boolean networks are needed to deepen our understanding of attractor dynamics under noise. It remains important to assess how generally these results hold, especially in large-scale networks where one can use sampling-based or approximate methods, which represents an important direction for future work. In addition, future research could explore how specific topological features, such as modularity [19, 47], redundancy [18, 48], and/or the presence of canalizing functions [49, 15, 16]. Extending the framework to include asynchronous updates [13], multi-node perturbations [36], or correlated noise [50] may provide further insight into biological function.
@@ Acknowledgements.
This research was supported in part by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. 2020R1I1A3068803), by Global - Learning & Academic research institution for Master’s $\cdot$ PhD students, and Postdocs (LAMP) Program of the National Research Foundation of Korea (NRF) grant funded by the Ministry of Education (No. RS-2024-00445180) and by the IITP (Institute of Information & Coummunications Technology Planning & Evaluation)-ITRC (Information Technology Research Center) grant funded by the Korea government (Ministry of Science and ICT) (IITP-2025-RS-2024-00437284). This work was conducted during the research year of Chungbuk National University in 2024-2025. The research of R.L. was partially supported by the grants NSF DMS-2424635, NIH R01 AI135128, and NIH R01 HL169974-01.
Appendix A Characteristics of real-world Boolean networks
In this appendix, we summarize the properties of a set of real-world Boolean networks used in our analysis, obtained from the Cell Collective database [46]. For each network, we report its number of nodes, mean in-degree, and mean bias of its Boolean functions in Table. 1.
Table 1: Real-world Boolean networks
Boolean networks bias degree size
Cortical Area Development 0.163 2.8 5
Cell Cycle Transcription by Coupled CDK 0.306 2.111 9
Mammalian Cell Cycle 2006 0.270 3.5 10
Predicting Variabilities in Cardiac Gene 0.303 2.533 15
Cardiac development 0.303 2.533 15
CD4 T cell dynamics – Workshop 0.079 1.875 16
Neurotransmitter Signaling Pathway 0.414 1.375 16
SKBR3 Breast Cell Line Short-term ErbB Network 0.084 2.563 16
BT474 Breast Cell Line Short-term ErbB Network 0.075 2.875 16
HCC1954 Breast Cell Line Short-term ErbB Network 0.087 2.875 16
CD4+ T Cell Differentiation and Plasticity 0.065 4.333 18
T-LGL Survival Network 2011 Reduced Network 0.25 2.389 18
Human Gonadal Sex Determination 0.289 4.158 19
Mammalian Cell Cycle 0.2422 2.55 20
B cell differentiation 0.296 1.773 22
Aurora Kinase A in Neuroblastoma 0.309 1.869 23
T cell differentiation 0.325 1.478 23